Please help me solve this equation:
3x-2y=5
x=y+10

Answer by jim_thompson5910(35256)

3x-2y=5

3(y+10)-2y=5

3y+30-2y=5

y+30=5

y=5-30

y=-25
--------------------------

x=y+10

x=-25+10

x=-15


So the solutions are x = -10 and y = -25
-------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim

-------------------------------------------------------------------------------------------------